Order of operations with equations
-4(2x – 3) = -8x + 5

Answer by checkley71(8403)

-4(2X-3)=-8X+5 OR -8X+12=-8X+5 OR 12=5 NO SOLUTION
ARE YOU SURE YOU HAVE THE CORRECT PROBLEM??????????????